### v1

# Hundred-Kilobyte Lookup Tables for Efficient Single-Image Super-Resolution

Abstract
Conventional super-resolution (SR) schemes make heavy use of convolutional neural networks (CNNs), which involve intensive multiply-accumulate (MAC) operations, and require specialized hardware such as graphics processing units. This contradicts the regime of edge AI that often runs on devices strained by power, computing, and storage resources. Such a challenge has motivated a series of lookup table (LUT)-based SR schemes that employ simple LUT readout and largely elude CNN computation. Nonetheless, the multi-megabyte LUTs in existing methods still prohibit on-chip storage and necessitate off-chip memory transport. This work tackles this storage hurdle and innovates hundred-kilobyte LUT (HKLUT) models amenable to on-chip cache. Utilizing an asymmetric two-branch multistage network coupled with a suite of specialized kernel patterns, HKLUT demonstrates an uncompromising performance and superior hardware efficiency over existing LUT schemes.
Introduction
Single image super-resolution (SISR) is a long-standing problem in computer vision that involves generating a high-resolution (HR) image from a low-resolution (LR) image. The goal is to recover high-frequency details that are lost in a downsampled image. Classical non-deep-learning methods can be divided into two main categories: interpolation and sparse coding. Interpolation approaches (Keys 1981; Kirkland and Kirkland 2010), such as bicubic interpolation, are simple and computationally efficient. They involve upsampling the LR image using a fixed kernel, e.g., a bicubic filter. However, they often suffer from limited generalizability and produce blurry images. Sparse-coding approaches (Timofte, De Smet, and Van Gool 2013, 2015), on the other hand, can achieve better results than the former but are computationally expensive and require hand-crafted features.
Figure 1: PSNR (dataset: Set5) vs. model size: HKLUTs are the first hundred-KB series of LUT-based SISR models targeting the sub-1MB regime (NB. $x$-axis is in log scale).
Like many other problems in computer vision, deep learning has revolutionized SISR. The mapping between LR and HR images can be learned directly from data, often resulting in super-resolved images of much better quality than classical methods. However, deep neural networks (DNNs) require a large number of floating-point operations, resulting in a high power budget. Recently, researchers have explored new schemes that combine lookup tables (LUT) with deep-learning SISR (Jo and Kim 2021; Li et al. 2022; Ma et al. 2022). Specifically, this approach pre-computes all possible input-output pairs of a deep SR network and stores them into a LUT for retrieval, thus skipping expensive computation during inference. As a result, LUT-based methods achieve inference speeds comparable to interpolation methods (Jo and Kim 2021), but produce results with a learning-based quality. Moreover, LUT-based approaches are agnostic to software and hardware platforms, making them much more versatile than DNNs which require dedicated software and hardware support. Nonetheless, recent LUT-based SR schemes (Li et al. 2022; Ma et al. 2022) mainly focus on improving performance by enlarging the receptive field (RF), while overlooking the memory constraints. This contradicts the trending edge AI that often runs on edge devices strained by limited resources, e.g., the popular Ultra96 FPGA board has $<1$MB on-chip memory. In fact, existing LUT-based schemes generally require multi-megabyte LUTs where off-chip storage is inevitable. In addition, existing LUT-based approaches may still require expensive interpolation and/or floating-point operations, resulting in higher peak memory and energy costs. This highlights the challenge as well as the need for significant LUT size reduction and better architecture design to address the storage and power issues, while upkeeping performance.
To this end, this work systematically explores the relationship between the number of input pixels, RF, and SR performance. Our study leads to the design of a family of models called Hundred-Kilobyte Look-up Tables (HKLUTs) which offer a compelling balance between memory and performance. To eliminate the time-consuming and energy-intensive interpolation, a two-branch architecture as in SPLUT (Ma et al. 2022) is employed. Yet instead of its symmetric branches (Ma et al. 2022), we exploit the idea of effective receptive field (ERF) (Luo et al. 2016) in different branches and reveal a larger RF is only required for the most-significant-bit (MSB) branch representing contextual semantic information, whereas a much smaller RF is sufficient for processing the least-significant-bit (LSB) branch.
This key insight leads to an innovative asymmetric parallel structure, which largely reduces the storage space by nearly half without sacrificing performance. Next, we extend our model to multistage by cascading LUTs to enlarge the RF. Our proposed architecture, in contrast to prior works, allows information exchange between two branches during stage transition, resulting in an uncompromising performance. Moreover, we are the first to utilize progressive upsampling on LUT-based approaches, which elegantly reduces the LUT size by half while even improving the final performance. Compared to other state-of-the-art (SOTA) LUT-based schemes, HKLUTs deliver competitive performance while being >10×>10$\times$> 10 × smaller in size (cf. Fig. 1). The main contributions of this paper are threefold:
  • We systematically study the relationship between RF, the number of input pixels, and performance in the SR task, leading to our design of extra lightweight LUTs.
  • We investigate the ERFs of two branches and propose an asymmetric parallel structure to reduce storage by 2$\approx$ 2$\times$ 2 × without sacrificing performance.
  • A novel multistage architecture is proposed which adopts a progressive upsampling strategy and enables the communication between the two branches during stage transition. This leads to an improved performance together with another 2×2$\times$2 × reduction in size.
By seamlessly integrating the above techniques, HKLUTs constitute the smallest LUT-based SISR schemes in the literature offering a highly competitive performance.
Related Work
The quest for high-quality SR on edge devices has fueled the interest and research of effective and efficient SISR schemes in recent years, which can be broadly categorized into two types: classical and deep-learning methods.
Classical Approaches. Basic linear interpolation methods (Keys 1981; Kirkland and Kirkland 2010), though simple and efficient, are limited in their abilities to adapt to the image content and often result in blurry images. Early example-based work (Timofte, De Smet, and Van Gool 2013, 2015) learned sparse dictionaries over pixels or patches to build a compact representation between LR and HR pairs. During inference, the HR output is obtained using the pre-computed projection matrix with the input features. With low computation requirements, RAISR (Romano, Isidoro, and Milanfar 2016) proposed an effective sharpening algorithm to enhance the visual quality of over-smoothed images caused by cheap interpolation methods. However, such sparse coding models remain time-consuming to compute the sparse representations, while RAISR still needs local gradient statistics and fails to provide optimal solutions due to its non-differentiable hashing operations.
Deep Learning-based Approaches. Deep learning methods have shown promising results in reconstructing image details for accurate super-resolution. Many deep learning SR works have been developed to reduce computational burdens by carefully adjusting the network structures, including, ESPCN (Shi et al. 2016a), CARN (Ahn, Kang, and Sohn 2018), RRDB (Wang et al. 2018) and IMDN (Hui et al. 2019). FALSR (Chu et al. 2021) proposed using neural architecture search (NAS) for both micro (block structure) and macro (layer-wise connection) levels. LapSRN (Lai et al. 2017) and its extension MS-LapSRN (Lai et al. 2018) employ Laplacian Pyramid to generate multi-scale predictions and progressively reconstruct HR images in multiple steps. FMEN (Du et al. 2022) shined in the NTIRE 2022 challenge with its low memory cost and short runtime through the enhanced residual block and high-frequency attention module. Although some models can be executed in real-time on GPUs, it is impractical to deploy them on edge devices with constrained hardware resources.
LUT-based Approaches. While deep learning-based methods can produce impressive results, they often require heavy computations and large storage. To accelerate the inference of DNNs, dedicated software and hardware platforms such as CUDA and GPUs are required. This poses challenges for cost-sensitive and resource-restrictive edge devices such as smartphones and TVs. To address this, SRLUT (Jo and Kim 2021) fuses DNN and LUT for SR applications by pre-computing and caching input-output pairs of a pre-trained SR network into a LUT for retrieval at test time. This approach eliminates expensive computation during DNN inference and achieves similar speed but better SR results than classical approaches. The number of input pixels is restricted to 4, preventing exponential growth in storage space for the resulting LUT. The rotation ensemble is proposed for both training and inference to improve SR performance. To further reduce the LUT size, input values are discretized into a fewer number of fixed levels, while interpolation is used to recover the quantization loss. Two independent works around the same time, namely, MuLUT and SPLUT (Li et al. 2022; Ma et al. 2022), propose to expand the RF covered by cascading LUTs. MuLUT proposes the use of multiple complementary LUTs together with the rotation ensemble to solve the intrinsic limitation of a small RF in SRLUT. It also designs a LUT-aware fine-tuning strategy to boost performance. However, MuLUT preserves the costly interpolation step of SRLUT. On the other hand, SPLUT (Ma et al. 2022) exploits a novel parallel framework where the quantization loss is compensated by using two parallel branches that separately handle the most significant 4 bits and least significant 4 bits. This eliminates the time-consuming interpolation step for a faster inference speed. SPLUT also suggests enlarging the RF through feature aggregation along the horizontal and vertical directions. Nonetheless, despite showing improved performance compared to SRLUT, both SPLUT and MuLUT require several times more storage space.
Development of HKLUT
Reduced Number of Input Pixels
The storage size of a LUT can be calculated using the formula $v^{n}\times r^{2}$, where $v$, $n$, and $r$ are the number of quantization levels, number of input pixels, and upscaling factor, respectively. The first term $v^{n}$ determines the number of entries, whereas the second term $r^{2}$ refers to the length of the output vector of each LUT. It is obvious that reducing the number of input pixels $n$ can result in an exponential reduction in storage space. Previous studies have primarily focused on expanding the RF by utilizing a fixed number of pixels (say, 4) to improve the final performance. SPLUT (Ma et al. 2022) uses feature aggregation to expand the RF, namely, the output feature maps generated from the previous LUT are divided into groups along the channel dimension and then padded horizontally or vertically. The padded feature maps are then added to generate a single-channel feature map, which serves as the input for subsequent LUTs. However, in order to generate multi-channel feature maps, the size of intermediate LUTs grows linearly with the number of output channels. This goes against our primary goal, which is to reduce storage space. Subsequently, we deliberately omit feature aggregation. On the other hand, SRLUT (Jo and Kim 2021) and MuLUT (Li et al. 2022) employ rotation ensemble to increase the area being covered. The final prediction $\hat{y}_{i}$ is computed as follows:
$\hat{y}_{i}=\frac{1}{N}\frac{1}{M_{k}}\sum_{k=0}^{N}\sum_{j=0}^{M_{k}}R^{-1}_{j}(LUT_{k}(R_{j}(x_{i}))$ (1)
where $x_{i}$ is the LR input, $LUT_{k}$ is the $k$th LUT, $R_{j}$ is the $j$th rotation operation, $N$ is the number of kernels (or LUTs), and $M_{k}$ is the number of rotation operations for the corresponding kernel. We remark that rotation ensemble, unlike feature aggregation, does not require extra storage.
Figure 2: Three types of kernels covering a $3\times 3$ RF with rotation ensemble via different numbers of input pixels. (a) SRLUT (Jo and Kim 2021) (b) HDLUT (c) LLUT.
Method RF # Pixels LUT size PSNR(dB)
SRNet $3\times 3$ 4 1.27 MB 29.87
LNet $3\times 3$ 3 76.77 KB 29.67
HDNet $3\times 3$ 2 9.03 KB 29.45
Table 1: Effect of the number of input pixels on PSNR, evaluated on Set5 dataset. “Net” is used to distinguish full-precision CNNs from quantized LUTs.
An important aspect that has been overlooked by previous works is that achieving the same RF is indeed feasible with fewer pixels. To fill in this gap, we design LLUT and HDLUT (named after their input-pixel shapes, cf. Fig. 2) to explore the relationship between the number of input pixels and the final PSNR subject to a prescribed RF size. Specifically, LLUT and HDLUT are designed to completely cover a $3\times 3$ area without any overlapping pixels, except for the pivot pixel, using a rotation ensemble. Table 1 shows the PSNR improves with an increasing number of input pixels, which is consistent with the intuition that having more pixels gathers more information, at the cost of a higher storage, viz. a 4-pixel LUT requires megabytes of storage, whereas the performance of the 2-pixel LUT degrades obviously. To balance storage and performance, we devise HDBLUT, which consists of three 3-pixel kernels. Using rotation ensemble, it covers an entire $5\times 5$ RF without any overlap, except for the pivot pixel (Fig. 3). In contrast to SRLUT and MuLUT, which have a $3\times 3$ RF with 4 overlapping pixels and a $5\times 5$ RF with 11 overlapping pixels, respectively, our designs, HDLUT and HDBLUT, aim to match the same RF with minimal input pixels. This achieves a near-optimal tradeoff between memory footprint and performance (a comparison of the RF of our designs vs others is in the Appendix). Using HDBLUT and HDLUT, along with the asymmetric parallel structure discussed in the sequel, we are able to outperform SRLUT with >6×>6$\times$> 6 × less storage.
Figure 3: With rotation ensemble, HDBLUT can cover a $5\times 5$ area with 3 three-pixel kernels. On the right, each square represents a single pixel, with a color indicating which kernel covers it.
Asymmetric Parallel Structure
Previous studies have attempted to reduce the size of LUTs by limiting the number of quantization levels to 17 and using interpolation to recover the quantization loss (Jo and Kim 2021; Li et al. 2022). However, the interpolation stage is time-consuming and energy-intensive, hindering the practicality on edge devices. To address this, SPLUT (Ma et al. 2022) proposes a parallel branch structure that avoids the need for interpolation. The structure divides the 8-bit input image into its most and least significant bits, viz. MSB and LSB branches each with 4 bits (viz. 16 levels) of information. The final SR image is obtained by fusing the results of the two branches at the last stage. While this approach gives promising results, it overlooks an important fact, namely, the effective receptive fields (ERFs) of MSB and LSB branches are completely different as they represent different kinds of information. Therefore, using the same sets of LUTs for both MSB and LSB branches might not be an optimal choice.
We trained two identical SRNets, using a $5\times 5$ kernel for both MSB and LSB branches. The results from the two branches are summed and gradients are computed using mean-squared error (MSE) loss. Fig. 4 shows the ERF, in terms of normalized gradients obtained from randomly selected $4\times 4$ regions of the SR result with respect to inputs, using the baby image from the Set5 dataset (cf. Fig. 5) (more demonstrations can be found in the Appendix). Although both MSB and LSB are processed under the same theoretical RF (i.e., $5\times 5$), the results show that the MSB branch, which captures the high-level contextual semantic information, has a larger ERF. On the other hand, a smaller ERF is sufficient for the LSB branch to describe fine details. This observation justifies an asymmetric parallel structure that exploits the differences between MSB and LSB to save storage space. By using a set of 3-pixel LUTs (i.e., HDBLUT) corresponding to CNNs with a $5\times 5$ RF for the MSB branch, and a set of 2-pixel LUTs (i.e., HDLUT) corresponding to CNNs with a $3\times 3$ RF for the LSB branch, we can nearly halve the storage space without compromising performance (cf. Table 3).
Figure 4: Four randomly selected effective receptive fields (ERFs). Top: Most significant 4 bits. Bottom: Least significant 4 bits. The brightness denotes model’s sensitivity to that pixel, justifying the assymetric $5\times 5$ and $3\times 3$ kernels for respective branches.
Figure 5: Overall architecture of the proposed method (HKLUT-S).
Multistage
LUT-based methods (Ma et al. 2022; Li et al. 2022) have demonstrated the efficacy of cascaded stages to enlarge the RF for better performance. SPLUT (Ma et al. 2022) splits the image into the MSB and LSB branches and deals with them separately with uniform quantization until the final addition, which prevents each branch from getting information from the other. MuLUT (Li et al. 2022) utilizes quantization and interpolation to build the connection between stages, which requires later fine-tuning to compensate for the loss of information caused by quantization.
Here, we propose a new framework to construct the cascaded stages as illustrated in Fig. 5. In each stage, we split the 8-bit input image into MSB and LSB branches for different LUT blocks to extract the features and upscale the resolution, then combine the features from both branches to compose a new enlarged image. The input and output of each stage are 8-bit images with different resolutions if the upscaling factor is not equal to 1. Such a module can then be stacked to constitute cascaded stages, wherein the latter stage can attain information from the two branches of the former stages and residual connections can be conveniently implemented. The entire flow requires no interpolation, which effectively speedups inference.
Furthermore, previous works utilize cascaded stages to increase the RF for a better performance, where upsampling is usually done in the last step. It is intuitively difficult to reproduce a high-resolution patch from a few pixels in one single step, especially when the upscaling factor is large. Besides, the storage space of LUT is proportional to the square of the upscaling factor $r$, we can save half the storage space by dividing the upscaling factor 4×4$\times$4 × (viz. $4\times 4=16$) into two factors 2×2$\times$2 × (viz. $2\times 2+2\times 2=8$). This simple yet previously unexplored progressive upsampling trick gives another booster to our proposed SISR scheme. Table 3 verifies the effectiveness of progressive upsampling in our LUT-based methods, along with further savings in storage.
Table 2: Asymmetric parallel structure: a smaller LSB-branch RF yields negligible performance drops but large storage savings.
Method RF # Pixels Storage Set 5 Set 14 BSDS100 Urban100 Manga109
MSB LSB MSB/LSB MSB/LSB PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM
HD HD $3\times 3$/$3\times 3$ 2/2 16KB 29.39/0.835 26.79/0.730 26.38/0.692 23.73/0.689 26.28/0.825
HD HDB $3\times 3$/$5\times 5$ 2/3 200KB 29.40/0.835 26.79/0.730 26.39/0.692 23.73/0.689 26.29/0.825
HDB HD $5\times 5$/$3\times 3$ 3/2 200KB 29.98/0.848 27.13/0.741 26.61/0.702 24.01/0.700 26.88/0.837
HDB HDB $5\times 5$/$5\times 5$ 3/3 384KB 29.99/0.849 27.15/0.742 26.63/0.703 24.02/0.701 26.90/0.839
# Stages Upscale Storage Runtime Set5 Set14 BSDS100 Urban100 Manga109
PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM
$1$ $4$ 200KB 108.8ms 29.97/0.84 27.13/0.741 26.61/0.702 24.01/0.700 26.86/0.837
$2$ $1\times 4$ 213KB 115.5ms 29.99/0.848 27.14/0.742 26.62/0.702 24.02/0.700 26.86/0.837
$2$ $2\times 2$ 100KB 201.7ms 30.35/0.859 27.39/0.748 26.73/0.706 24.23/0.711 27.38/0.852
$3$ $1\times 2\times 2$ 112.5KB 212.1ms 30.34/0.860 27.40/0.748 26.75/0.707 24.25/0.712 27.46/0.853
$3$ $2\times 1\times 2$ 112.5KB 235.3ms 30.41/0.860 27.44/0.749 26.78/0.707 24.27/0.713 27.51/0.854
$3$ $2\times 2\times 1$ 112.5KB 475.0ms 30.43/0.861 27.47/0.750 26.79/0.707 24.31/0.713 27.58/0.856
Table 2: Asymmetric parallel structure: a smaller LSB-branch RF yields negligible performance drops but large storage savings.
Table 3: Ablating various progressive upsampling strategies. Runtime is measured for outputting a $1280\times 720$ image on a desktop CPU averaged across 10 runs. We refer to models in the lime and blue rows as HKLUT-S and HKLUT-L, respectively.
Method Size Ratio Set5 Set14 BSDS100 Urban100 Manga109
PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM
Nearest (Bevilacqua et al. 2012) - - 26.25/0.737 24.65/0.653 25.03/0.629 22.17/0.615 23.45/0.741
Bilinear (Kirkland and Kirkland 2010) - - 27.55/0.788 25.42/0.679 25.54/0.646 22.69/0.635 24.21/0.767
Bicubic (Keys 1981) - - 28.42/0.810 26.00/0.702 25.96/0.667 23.14/0.657 24.91/0.787
ANR (Timofte, De Smet, and Van Gool 2013) 1.43MB 14.64 29.70/0.842 26.86/0.737 26.52/0.699 23.89/0.696 26.18/0.821
A+ (Timofte, De Smet, and Van Gool 2015) 15.17MB 155.34 30.27/0.860 27.30/0.750 26.73/0.709 24.33/0.719 26.91/0.848
CARN-M (Ahn, Kang, and Sohn 2018) 1.59MB 16.28 31.82/0.890 28.29/0.775 27.42/0.731 25.62/0.769 29.85/0.899
RRDB (Wang et al. 2018) 63.83MB 653.62 32.60/0.900 28.88/0.790 27.76/0.743 26.73/0.807 31.16/0.916
SRLUT (Jo and Kim 2021) 1.27MB 13.05 29.82/0.847 27.01/0.736 26.53/0.695 24.02/0.699 26.80/0.838
SPLUT-S (Ma et al. 2022) 5.5MB 56.32 30.01/0.852 27.20/0.743 26.68/0.702 24.13/0.706 27.00 0.843
SPLUT-M (Ma et al. 2022) 7MB 71.68 30.23/0.857 27.32/0.746 26.74/0.704 24.21/0.709 27.20 0.848
SPLUT-L (Ma et al. 2022) 18MB 184.32 30.52/0.863 27.54.0.752 26.87/0.709 24.46/0.719 27.70 0.858
MuLUT-SDY (Li et al. 2022) 3.82MB 39.15 30.40/0.860 27.48/0.751 26.79/0.709 24.31/0.714 27.52/0.855
MuLUT-SDY-X2 (Li et al. 2022) 4.06MB 41.60 30.60/0.865 27.60/0.754 26.86/0.711 24.46/0.719 27.90/0.863
HKLUT-S (ours) 100 KB 1 30.35/0.859 27.39/0.748 26.73/0.706 24.23/0.711 27.38/0.852
HKLUT-L (ours) 112.5 KB 1.13 30.41/0.860 27.44/0.749 26.78/0.707 24.27/0.713 27.51/0.854
Table 4: Different SR methods on 5 benchmark datasets. The size ratio is w.r.t. HKLUT-S. The 4 groups from top to bottom are interpolation, sparse-coding, CNN & LUT approaches, respectively.
Experiments
Implementation Details
Datasets and Metrics. We train the proposed HKLUT on the DIV2K dataset (Agustsson and Timofte 2017), a popular dataset in the SR field. The DIV2K dataset provides 800 training images and 100 validation images with 2K resolution and the corresponding downsampled images. We use the widely adopted Peak Signal-to-Noise Ratio (PSNR) and structural similarity index (SSIM) (Wang et al. 2004) as evaluation metrics. Five well-known datasets: Set5 (Bevilacqua et al. 2012), Set14 (Zeyde, Elad, and Protter 2012), BSDS100 (Martin et al. 2001), Urban100 (Huang, Singh, and Ahuja 2015) and Manga109 (Matsui et al. 2017) are benchmarked. Following other LUT-based methods, we mainly focus on SR tasks with an upscaling factor of 4 (i.e., 4×4$\times$4 × SR). Besides, like MuLUT (Li et al. 2022), we measure and report the runtime, theoretical energy cost, and peak memory of generating a $1280\times 720$ HR image from a $640\times 360$ LR image.
Experimental setting. Before converting into LUTs, the network is trained with 800 training images in DIV2K for 200k iterations with a batch size of 16 on Nvidia RTX 3090 GPUs, each with a 24GB buffer. We utilize the Adam optimizer (Kingma and Ba 2014) ($\beta_{1}=0.9$, $\beta_{2}=0.999$ and $\epsilon=1\times 10^{-8}$) with the MSE loss to train the HKLUT. The initial learning rate is set to $5\times 10^{-4}$, which decays to one-tenth after 100k and 150k iterations, respectively. We randomly crop LR images into $48\times 48$ patches as input and enhance the dataset by random rotation and flipping. The runtime is measured on an Intel Core i5-10505 CPU with 16GB RAM, averaged over 10 runs.
Baselines. We evaluate HKLUT against several well-recognized SISR methods, such as interpolation-based methods (nearest neighbor (Bevilacqua et al. 2012), bilinear (Kirkland and Kirkland 2010) and bicubic (Keys 1981) interpolation), sparse-coding-based methods (ANR (Timofte, De Smet, and Van Gool 2013) and A+ (Timofte, De Smet, and Van Gool 2015)), DNN-based methods (CARN (Du et al. 2022) and RRDB (Wang et al. 2018)), and LUT-based methods (SRLUT (Jo and Kim 2021), SPLUT (Ma et al. 2022), and MuLUT (Li et al. 2022)).
Quantitative Results
Asymmetric Parallel Structure. To demonstrate the effectiveness of our proposed asymmetric parallel structure, we utilize the 2-pixel HDLUT with $3\times 3$ RF and 3-pixel HDBLUT with $5\times 5$ RF to investigate the relationship between RF, LSB, and MSB for the 4×4$\times$4 × SR. Table 3 shows that using HDLUTs for both branches results in worse performance (29.39dB for Set5). A negligible increase (+0.01dB) in performance is achieved when we replace the LSB branch with the larger-RF HDBLUT, at a cost of over 12 times the growth in model size (200KB vs. 16KB). However, by simply swapping the two branches, i.e., using HDBLUT for MSB and HDLUT for LSB, we obtain a significant boost in performance (+0.58dB) with the same storage and computing. When using HDBLUT for both MSB and LSB branches, we only obtain a marginal increase in PSNR (+0.01dB), at a cost of nearly doubling the storage (384KB vs. 200KB). This validates our conjecture that a 2-pixel LUT is sufficient for the LSB because of the small ERF, while a 3-pixel LUT is necessary for the MSB to capture contextual information such as texture. In short, the proposed asymmetric parallel structure allows for the efficient allocation of resources and saves half of the storage space without sacrificing performance.
Multistage. Table 3 examines various upsampling strategies to demonstrate the effectiveness of progressive upsampling. Following the asymmetric parallel structure, HDBLUT and HDLUT are used for each stage’s MSB and LSB branches, respectively. Comparing the 1st and 3rd rows, using an asymmetric parallel structure, 2-stage progressive upsampling ($2\times 2$) shows a clear advantage in terms of SR performance compared to 1-stage upsampling (4×4$\times$4 ×) (30.35dB vs. 29.97dB on Set 5), while requiring only half the storage space (100KB vs. 200KB). The advantage of progressive upsampling becomes even clearer when compared to results on the 2nd row, which corresponds to the same 2-stage but with upsampling in a single step ($1\times 4$) (-0.36dB). Our results show that $2\times 1\times 2$ upsampling achieves the best tradeoff by considering storage, runtime, and performance. We, therefore, refer to models listed on the 3rd and 5th rows as HKLUT-S (highlighted in lime) and HKLUT-L (highlighted in blue), respectively.
Performance and Storage In Table 4, we compared our proposed HKLUT with well-recognized SISR approaches such as interpolation, sparse-coding, and CNNs, as well as the latest LUT-based methods. Our approach has a clear advantage in storage, which aligns with the objective of this paper. HKLUT-S, being the smallest parameterized model, only requires 100KB, which is 10×10$\times$10 × less than the second smallest SRLUT (1.27MB) with a much better performance. Both HKLUT-S and HKLUT-L outperform all classical baselines, including A+, which is significantly larger than HKLUT-S (15.17MB vs. 100KB). Although CNN-based approaches exhibit good performance in terms of PSNR and SSIM, they often require larger storage. For example, RRDB is 650×650$\times$650 × larger than our HKLUT-S, making it difficult to deploy on edge devices. With respect to LUT-based approaches, our approaches not only take up the least storage, but they also upkeep comparable performance. HKLUT-L outperforms MuLUT-SDY on the Set5 dataset (30.41dB vs. 30.40dB) with 35×35$\times$35 × less storage. Despite being 70×70$\times$70 × smaller, HKLUT-S outperforms SPLUT-M by a significant margin (30.35dB vs. 30.23dB on Set5).
Figure 6: Energy cost and peak memory to output a $1280\times 720$ image through 2×2$\times$2 × SR.
Energy & Peak Memory. Following the setting in MuLUT (Li et al. 2022), Fig. 6 shows the theoretical energy costs and peak memory usages of FSRCNN (32.70dB) and several LUT-based methods (31.73dBsimilar-$\sim$32.35dB) for generating a $1280\times 720$ image through 2×2$\times$2 × SR. We estimated the theoretical energy cost using the same protocol as AdderSR (Song et al. 2021), and computed the peak memory usage using memray [^1] Measured with memray (https://github.com/bloomberg/memray) on a desktop CPU.. For LUT-based approaches, we used their smallest variants. One of the major advantages of LUT-based methods is their ability to drastically reduce energy and memory consumption. Even FSRCNN, the simplest CNN for SISR, consumes 900×900$\times$900 × and 7.5×7.5$\times$7.5 × more energy and memory compared to HKLUT. The feature map aggregation module of SPLUT not only instantiates larger LUTs but also requires floating-point operations, resulting in increased energy consumption and peak memory usage. Meanwhile, both SRLUT and MuLUT consume a significant amount of energy for the interpolation operation. In contrast, HKLUT gets rid of energy-intensive interpolation and solely relies on integer operations. Fig. 6 shows that HKLUT exhibits the lowest energy cost and peak memory usage amongst all competitors.
Figure 7: Qualitative results on selected images from five benchmark datasets.
Runtime. Using the same settings as for calculating energy and peak memory, we compared the runtimes of FSRCNN and LUT-based approaches on both a desktop CPU and the Raspberry Pi 4 Model B. To ensure a fair comparison, we used the official Python code released by the authors (Jo and Kim 2021; Li et al. 2022; Ma et al. 2022), without relying on native mobile implementation or any speedup tricks. We recorded the average runtime of generating a $1280\times 720$ image from a $640\times 360$ image in Table 5. It is clear that the 4-simplex interpolation bottlenecks the inference speed of SRLUT and MuLUT at the final stage. SPLUT and our approaches, featuring parallel branches, exhibit shorter inference times on Raspberry Pi and CPU. Unsurprisingly, HKLUT achieves the fastest inference speed.
Method FSRCNN SRLUT SPLUT-S MuLUT-SDY HKLUT
CPU (ms) 332.64 1412.24 197.49 5128.17 148.67
Raspberry (ms) 3589.86 8021.56 1096.73 35637.12 908.52
Table 5: Comparing runtimes for generating a $1280\times 720$ image through 2×2$\times$2 × SR. Results are obtained by averaging across 10 runs.
Qualitative Results
Fig. 7 presents a visual comparison between the bicubic interpolation baseline and different LUT-based approaches on images selected from five benchmark datasets. The results clearly demonstrate that LUT-based approaches improve sharpness compared to bicubic interpolation which generates blurry images. The 5th row of images (Manga109) shows that the SRLUT approach exhibits severe rippling artifacts around the edges, which may be caused by its limited RF. HKLUT-S produces similar sharpness and clarity to MuLUT-SDY-X2 and SPLUT-L, despite being more than 40×40$\times$40 × and 180×180$\times$180 × smaller, respectively. Although A+ brings fewer artifacts, it is computation-heavy and not able to show clearer textures. Overall, the results show that LUT-based approaches outperform bicubic interpolation qualitatively, and our proposed HKLUT has been proven to be a highly competitive alternative to its larger-size LUT-based counterparts, yet at a drastically reduced storage cost.
Conclusion
This paper targets minimalist LUTs for efficient LUT-based SISR. We propose a new class of models called HKLUTs employing kernels with fewer input pixels while retaining the RF, resulting in an exponential reduction in storage without compromising performance. In particular, HKLUT consists of an asymmetric parallel structure that employs big-little kernel patterns to adapt to the underlying characteristics of MSB and LSB branches. A progressive multistage architecture, without expensive interpolation between stages, is devised to reduce storage by half. By seamlessly fusing these innovations, HKLUT models set the first record to attain $>30$dB on the Set 5 dataset at merely one-tenth of 1MB storage. Extensive experiments verify that HKLUT is a versatile alternative to previous large LUT-based approaches on resource-constrained devices.
Appendix
Details of Neural Nets during Training
During training, each look-up table (LUT) is replaced with a simple convolutional neural network (CNN), as shown in Fig. 8. The CNN consists of one convolution (CONV) layer with a $1\times K$ kernel, and five $1\times 1$ CONV layers, interleaved with RELU (Glorot, Bordes, and Bengio 2011) nonlinearity. No dense and residual connections are used here. $K$, $C$, and $S$ represent the number of input pixels to the LUT, the output channel number of the CONV layer (i.e., 64), and the upscaling factor, respectively. Since the LUT only accepts a finite number (K) of input pixels, all convolution kernels are of size $1\times 1$, except for the first one, which is of size $1\times K$. If the kernel shape of the first convolution is neither horizontal nor vertical, the pixels are first sampled according to the defined index pattern and then reshaped to $1\times K$. After the last CONV layer, pixel shuffle (Shi et al. 2016b) is applied to generate the high-resolution (HR) image. A Tanh activation function is used at the end to project the output to a range between -1 and 1. To cache CNN outputs into a LUT, they are converted from FP32 to INT8. This is done by multiplying the FP32 values by 127 to ensure they fit within the INT8 range (-128 to 127). The resulting numbers are rounded down to the nearest integer and stored in the LUT as INT8.
Figure 8: Detailed architecture of the neural network employed during training.
(a) Set14
(b) B100
(c) Urban100
(d) Manga109
Figure 9: Four randomly selected ERFs obtained from images in the Set14, B100, Urban100, and Manga109 datasets. Top: MSB. Bottom: LSB. The brightness denotes the model’s sensitivity to that pixel.
Effective Receptive Field
As elaborated in the manuscript, SPLUT (Ma et al. 2022) neglects the disparity between the information contained in the most-significant-bit (MSB) and the least-significant-bit (LSB) branches, despite their identical theoretical receptive fields (e.g., $5\times 5$). We hypothesize that using the same sets of LUTs for both MSB and LSB branches may not be the optimal choice. To find an ideal solution, we visualize the effective receptive fields (ERFs) of both branches and introduce an asymmetric structure based on this observation, aiming to substantially reduce the size of LUTs. In Fig. 2, we provide additional demonstrations of the ERFs of various images across datasets, which confirm that the MSB branch has a larger ERF while the LSB branch maintains a more compact one.
LUT Design Comparison
In this section, we highlight the differences in LUT designs used in various LUT-based SISR methods. Fig. 3 shows a comparison of different LUT designs. Unlike SRLUT and MuLUT, which use a $3\times 3$ RF with 4 overlapping pixels and a $5\times 5$ RF with 12 overlapping pixels, respectively, our designs, HDLUT and HDBLUT, aim to match the same RF with minimal input pixels. This achieves a near-optimal tradeoff between memory footprint and performance.
(a) SRLUT
(b) HDLUT
(c) MuLUT
(d) HDBLUT
Figure 10: A comparison of various LUTs’ RF patterns. The numbers indicate all overlapping pixels, with darker pixels showing locations where the same kernel overlaps with itself during rotation ensemble, while pixels with two distinct colors denote areas where different kernels overlap with each other. In each case, the center pixel is used as a pivot for rotation by all kernels.
Architectural Design Comparison
This section compares the architectural designs of different LUT-based approaches for SISR. Fig. 11 shows block diagrams that illustrate the overall architectures of these approaches. In this context, LUT refers to a CNN with a structure similar to that of the CNN shown in Fig. 1. The prefix or subscript describes the kernel shape of the first CONV layer of this CNN.
Fig. 11 (a) depicts the overall structure of SRLUT (Jo and Kim 2021). SRLUT consists of only a single LUT, which corresponds to a CNN with the first layer using a $2\times 2$ square kernel. During both training and inference, rotation ensemble is used to expand the receptive field (RF) from $2\times 2$ to $3\times 3$. MuLUT (Li et al. 2022), as illustrated in Fig. 4 (c), extends SRLUT by using multiple LUTs with different input kernel shapes in parallel. These kernels have complementary indexing patterns, as they can fill up the entire $5\times 5$ area using the rotation ensemble trick. The same structure can be stacked in a sequential manner to form a multistage structure to further enlarge the RF.
Fig. 4 (b) shows the overall structure of SPLUT (Ma et al. 2022). SPLUT first divides an 8-bit image into two branches: one with the most significant 4 bits (MSB) and the other with the least significant 4 bits (LSB) to eliminate interpolation at the final stage. To maintain consistency, we follow the terminology from SPLUT (Ma et al. 2022). Here, $\rm{LUT}_{HW}$ is referred to as spatial lookup, which uses the $2\times 2$ square kernel as the S-LUT in Fig. 4 (a) and (c), but with multi-channel outputs. The number of output channels can be either 4 (SPLUT-S), 8 (SPLUT-M), or 16 (SPLUT-L). The multi-channel feature maps are then grouped along the channel dimension and fed to either vertical or horizontal aggregation. The aggregated features are then fed to $\rm{LUT}_{HC}$ and $\rm{LUT}_{WC}$, both of which have a $2\times 2$ input pattern. The input pattern of $\rm{LUT}_{HC}$ is along both the height and channel dimensions, while that of $\rm{LUT}_{WC}$ is along both the width and channel dimensions. The aggregation modules, $\rm{LUT}_{WC}$ and $\rm{LUT}_{HC}$, can be repeated multiple times in series to further expand the RF. The final super-resolved (SR) image is produced by summing the outputs of the two branches.
Figure 11: Comparison of architectural designs. Residual connections are omitted for brevity. (a) SRLUT (Jo and Kim 2021) (b) SPLUT (Ma et al. 2022) (c) MuLUT (Li et al. 2022) (d) HKLUT (Ours)
The proposed HKLUT architecture is illustrated in Fig. 4 (d). Our method utilizes the two-branch structure of SPLUT to eliminate the interpolation step. However, instead of employing the same set of LUTs in both branches, we propose an asymmetric parallel structure that uses smaller LUTs for the LSB branch to reduce storage space. Within each branch, we use multiple LUTs with the rotation ensemble trick similar to MuLUT. However, LUTs of our design have fewer input pixels, resulting in an exponential reduction in size. Summing the results from the two branches yields the final SR image. The entire structure is treated as a building block. A multistage structure is constructed by repeating the whole module. This allows convenient implementation of different progressive upsample strategies to further enhance the results and reduce storage.
Ablation Studies
Ablation on LUT Cascading
In previous work (Ma et al. 2022), a series-parallel structure is proposed to eliminate the need for interpolation. Such parallel structure divides an 8-bit image into MSB and LSB parts to be handled separately, while the series structure refers to cascading LUTs within individual branches. Communication between two branches is only allowed at the end of the model to generate the final SR image by merging information from two branches. The output activations of each LUT in intermediate stages are clipped between [-8, 7] and quantized into 16 levels. However, we find that this strategy is counter-intuitive and may not be optimal. On the contrary, our way of cascading LUTs allows communication between the two branches during stage transition. The output of each stage is clipped between [0, 255] represented in INT8, which allows the generation of meaningful results even in intermediate stages. Table 6 shows that our approach significantly outperforms the multistage structure in SPLUT (Ma et al. 2022) (30.35dB vs. 29.32dB for Set5), both adopting the $2\times 2$ progressive upsampling technique. Note that using the SPLUT way of cascading LUTs without feature aggregation performs even worse than our single-stage method (cf. 1st row of Table 3 in the manuscript).
Ablation on Residual Connection
For computer vision tasks, residual connection (He et al. 2016) is widely used to facilitate the training of neural networks. Various papers (Lai et al. 2017; Ahn, Kang, and Sohn 2018; Ma et al. 2022) have proven the effectiveness of residual connection to maintain the feature precision and improve the SR performance. By using residual connections, the model focuses on reconstructing the residual information between the ground truth HR image and the upsampled image, which intuitively reduces the difficulty of the underlying SR task. Table 7 compares the results of the three commonly used upsampling methods (nearest, bilinear, and bicubic). The effects of the three methods are indistinguishable. To reduce runtime, we selected the nearest interpolation as our upsampling strategy for residual connections.
Method Set5 Set14 BSDS100
PSNR/SSIM PSNR/SSIM PSNR/SSIM
HKLUT-S 30.35/0.859 27.39/0.748 26.73/0.706
SPLUT 29.32/0.836 26.55/0.724 26.15/0.686
Table 6: Ablation on different ways of cascading LUTs. HDBLUT and HDLUT are used for MSB and LSB branch, respectively.
Method Set5 Set14 BSDS100
PSNR/SSIM PSNR/SSIM PSNR/SSIM
Nearest 30.35/0.859 27.39/0.748 26.73/0.706
Bilinear 30.35/0.859 27.39/0.748 26.75/0.706
Bicubic 30.36/0.859 27.39/0.749 26.75/0.706
Table 7: Ablation on different ways of implementing residual connection. “Nearest” is the same as HKLUT-S.